OpenStudy (anonymous):

x + y = 6 x - y = 8 Solve the system of equations.

OpenStudy (anonymous):

@TuringTest

OpenStudy (turingtest):

ok, well you can add those two equations together do you know how to do that?

OpenStudy (anonymous):

no /: I'm in oddesy and don't have any help

OpenStudy (jusaquikie):

solve both for Y algebraically. ie Y= 6-x set both equal ie (6-X) = (???) Solve for X plug in your value for X into one of the original equations and solve for Y

OpenStudy (turingtest):

or you can do the following\[~~~~~x+y=6\\+(x-y=8)\]just add the terms vertically

OpenStudy (anonymous):

so it would be (7,-1)?

OpenStudy (texaschic101):

you can check your answers by subbing them back into the equations to see if they are true

OpenStudy (turingtest):



OpenStudy (turingtest):

let's just look at the x termswe are adding x+x what is that?